OpenStudy (konradzuse):

question about fractional division i.e., 1/(1/3)

OpenStudy (konradzuse):

Originally I thought that when you get something over a fraction you want to make it into fraction * fraction i.e, in this case 1/1 * 1/3?

OpenStudy (konradzuse):

Apparently in this case the answer is 3, so it's goign to be 1 * 3/1 = 3?

OpenStudy (phi):

multiply top and bottom by 3 3/3

OpenStudy (konradzuse):

hmm?

OpenStudy (phi):

\[ \frac{3}{\frac{1}{3}}\cdot \frac{3}{3} \]

OpenStudy (konradzuse):

oh ok.

OpenStudy (phi):

oops make that a 1 up top in the first fraction

OpenStudy (konradzuse):

shouldn't it be 1/(1/3)

OpenStudy (konradzuse):

kk

OpenStudy (konradzuse):

Never have seen this method, like I said normally I just flip and multiply, but I always though the num/dom was one and the next ones were just multi, but flipped.

OpenStudy (phi):

flip and multiply is "short cut" to doing it the way I showed.

OpenStudy (konradzuse):

I gues smaybe because they were in paren? I feel like I've done it the other way before and it worked, so maybe I'm just thinking of that way, but continue on please.

OpenStudy (anonymous):

\(\frac{1}{\frac{1}{3}}\) in english "the reciprocal of one third" or think "how many thirds are in one?"

OpenStudy (konradzuse):

:O that makes some sense there Satellite :).

OpenStudy (konradzuse):

but is that situational? On wolfram if I do 1/1/3 I get 1/3 but if I do 1/(1/3) I get 3?

OpenStudy (phi):

thinking about is always a good option. and stick to basic principles (multiply by 1 for example) rather than tricks, if you have any doubts

OpenStudy (phi):

On wolfram if I do 1/1/3 that is order of operations

OpenStudy (konradzuse):

but your method makes a lot of sense too, try to cancel out so in this case 1/3 * 3/3 = 3/(3/3) = 3/1 = 3 :)

OpenStudy (konradzuse):

oh okay.. stoopid wolfram :P.

OpenStudy (anonymous):

\[\frac{a}{b}/\frac{c}{d}=\frac{a }{b }*\frac{d}{c} \]

OpenStudy (konradzuse):

So no matter what it will always be 3 huh... grr...

OpenStudy (phi):

1/1/3 is being interpreted as \[ \frac{\frac{1}{1}}{3} \]

OpenStudy (konradzuse):

yuh phi :)

OpenStudy (konradzuse):

I knew that robto. in this case that would be 1/1/3/1 which would be 1/1 * 1/3 = 1/3...

OpenStudy (amistre64):

unless there are grouping symbols, the math is read right to left. 1/1/3 is read as: 1/1 = 1, 1/3 = 1/3 1/(1/3) tells us that (1/3) is evaluated first to get a different result

OpenStudy (amistre64):

pfft ... cant even get my directions correct lol; read left to right

OpenStudy (konradzuse):

:)